OpenStudy (anonymous):

Find the exact value of sin55°cos25°+cos55°sin25°

OpenStudy (welshfella):

I have a feeling the 25 may be 35 ??? WE can find and exact value then

OpenStudy (anonymous):

@jim_thompson5910 any ideas?

jimthompson5910 (jim_thompson5910):

sorry I don't know all I know is that sin55°cos25°+cos55°sin25° = sin(80°) I used the identity sin(x+y) = sin(x)*cos(y)+cos(x)*sin(y)

jimthompson5910 (jim_thompson5910):

I don't know how to find the exact value of sin(80°) in terms of radicals

jimthompson5910 (jim_thompson5910):

@welshfella has a good point, if the 25 was 35, then sin55°cos35°+cos55°sin35° = sin(55+35) = sin(90°) = 1